OpenStudy (anonymous):

How would I simplify: (sqrt 6w^3) * (12w)

OpenStudy (anonymous):

well your problem is \[\sqrt{6w^{3}}(12w)\] or ..... \[(6w^{3})^{\frac{1}{2}}(12w)\]

OpenStudy (anonymous):

soo you get... \[6^{\frac{1}{2}}(w^{\frac{3}{2}})(12)(w)\]

OpenStudy (anonymous):

and when you multiply you just add the exponents...

OpenStudy (anonymous):

Thanks!